A girl is half the age of her mum and her dad is 5 years ahead of her mum. If you add up their ages you get 100. What is the age of each one of them?
Answer by JLJL(8)

Let the girl's age = X
Let the mum's age = Y
Let the dad's age = Z
If the girl is half the age of her mum, then X = 1/2Y
If the dad is 5 years ahead of the mum, then Z = Y + 5
If the sum of age of all three people is 100, then X + Y + Z = 100
We can now replace X and Y in terms of Z.
Thus,
X + Y + Z = 100
1/2Y + Y + (Y + 5) = 100
1/2 Y + 2Y + 5 = 100
5/2Y + 5 = 100
5Y + 10 = 200
5Y = 200 - 10
5Y = 190
Y = 190/5
Y = 38
Substitute Y = 38 into X, then X = 1/2(38) = 19
Substitute Y = 38 into Z, then Z = 38 + 5 = 43
Conclusion: The girl's age is 19, the mum's age is 38 whereas the dad's age is 43 respectively.